# Dilaton-induced variations in Planck constant and speed of light

### v1

@@ Abstract
We highlight a hidden aspect of scale-invariant actions that allow matter to couple with a dilaton field:   the dynamics of the dilaton can be used to induce variations in Planck’s quantum of action $\hbar$ and the speed of light $c$. This mechanism for generating variable $\hbar$ and $c$ in curved spacetimes offers novel insights into the origin of late-time cosmic acceleration, bypassing the need for dark energy.
Motivation—Let us revisit the quantum electrodynamics (QED) of a spinor field $\psi$ with inertial mass $m$ and charge $e$, coupled with a $U(1)$ gauge vector field $A_{\mu}$ and embedded in a 4-dimensional spacetime described by the Einstein–Hilbert (EH) action of General Relativity (GR). With the participation of $\hbar$ and $c$ explicitly restored, the full action reads:
@@
However, this relationship between $G$ and $\phi$ could cease to hold if matter couples non-minimally with gravity via the BD field. In this scenario, the involvement of $\phi$ in the matter Lagrangian, in principle, can cause both $c$ and $\hbar$ to vary alongside $\phi$ in spacetime, representing a departure from the standard BD paradigm. The current paper explores this potentially far-reaching scenario.
@@
In the literature, when scale symmetry, also known as dilatation symmetry, is broken—specifically, when $\chi$ spontaneously acquires a non-vanishing vacuum expectation value—the process gives rise to a Goldstone mode conventionally referred to as a dilaton. We will adopt the terminology of “dilaton” for $\chi$ in this paper. We should note, however, that the field $\chi$ in our action $\mathcal{S}_{2}$ can vary in spacetime, hence forfeiting its translational invariance.
Interestingly, a scale-invariant action of gravity and matter, such as $\mathcal{S}_{2}$, is able to evade observational bounds on the fifth force, a result established in [2, 3]. One can also add a “potential” term $V(\chi)$ to $\mathcal{L}_{\chi}$ provided that the added term respects scale symmetry. For example, $V(\chi)$ may contain terms like $\chi^{4}$ and $\chi^{-4}(\nabla\chi)^{4}$. Nevertheless, the inclusion of a scale-invariant potential will not affect the consideration presented in the rest of this paper.
@@
$G=\frac{\hat{c}^{3}}{16\pi\hat{\chi}^{2}\hat{\hbar}}\,.$ (15)
Importantly, per Eq. (14), the charge $e$ and inertial mass $m$ are independent of $\chi$. This result is desirable, as both charge and inertial mass—being intrinsic properties of a particle—should be oblivious to external factors, in particular the background dilaton field $\chi$. Befittingly, in our scheme, rather than making mass variable, the effect of the dilaton is translated into $\hbar_{\chi}$ and $c_{\chi}$, which respectively regulate the quantization and propagation of fields against a background spacetime. [Note: It is essential to note that the dimensionless parameters $\alpha$ and $\mu$ are “running coupling constants” in the renormalization group flow of $\mathcal{L}_{\text{QED}}^{(\chi)}$ when loop corrections involving $\psi$ and $A_{\mu}$ are included. Thus, although $e$ and $m$ are independent of $\chi$, they can “run” as functions of the momentum scale at which they are measured.]
@@
Moreover, our scheme brings about an additional benefit: not only is the $U(1)$ gauge charge $e$—which specifies the strength of electromagnetic interaction—independent of $\chi$, but the Newton gravitational constant $G$—characterizing the strength of gravitational interaction—is also independent of $\chi$, as evident in Eq. (15). Note that in our scheme, the Planck mass, given by $m_{\chi}^{\text{Planck}}\,:=$$\sqrt{\frac{\hbar_{\chi}c_{\chi}}{G}}=$$\,\sqrt{16\pi}\,\frac{\hat{\chi}\hat{\hbar}}{\hat{c}}$, is constant rather than variable.  [^1] Our paper does not concern with the hierarchy issue, namely why $m_{\chi}^{\text{Planck}}\gg m$, or equivalently why $\mu\ll\sqrt{16\pi}$.
@@
Within the open set, $\mathcal{L}_{\text{QED}}^{\text{ eff}}$ describes the quantum electrodynamics of a spinor field with charge $e$ and inertial mass $m$, specified in Eq. (14), coupled with a $U(1)$ gauge vector field. It is important to emphasize that, unlike the original $\mathcal{L}_{\text{QED}}$ in Eq. (3), the quantization and propagation of fields in $\mathcal{L}_{\text{QED}}^{\text{ eff}}$ are governed by the effective quantum of action $\hbar_{\chi}$ and speed of light $c_{\chi}$, which in turn are determined by the dilaton $\chi$ via Eq. (13). More generally, as $\chi$ varies across the manifold, different locations then correspond to separate replicas of the effective Lagrangian $\mathcal{L}_{\text{QED}}^{\text{ eff}}$, each operating with its respective values of $\hbar_{\chi}$ and $c_{\chi}$. Conceptually, the dynamics of $\chi$ thus induces variations in $\hbar$ and $c$ across spacetime.
@@
$i\hbar_{\chi}\,\frac{\partial}{\partial t}\bigl{|}\Psi\bigr{\rangle}=\hat{H}\bigl{|}\Psi\bigr{\rangle}\,.$ (17)
A physical consequence is that, for two locations with different dilaton values $\chi_{1}$ and $\chi_{2}$, the evolution of a quantum state occurs at different rates determined by each respective $h_{\chi}$ value. The discrepancy in the clock rates at the two locations gives rise to a new time dilation effect, resulting from the variation in $\hbar$ induced by the dynamics of $\chi$. We will briefly discuss this concrete definitive prediction in a later section of the paper.
@@
The energy level of an (relativistic) electron in a quantum state $\left|n,\,j\right\rangle$ of a Hydrogen atom is a well-established result; it is given by [16]
@@
$A=\frac{4\alpha}{3}\frac{m^{3}c_{\chi}^{4}}{h_{\chi}^{3}}\left\langle\vec{r}_{if}\right\rangle^{2}\,.$ (26)
@@
This result is one of the key findings of our paper.  [^2] There are two examples in classical physics that carry a $3/2$–exponent hallmark: (1) Kepler’s Third law, which states that the square of a planet’s orbital period $T$ is proportional to the cube of the semi-major axis $a$ of its orbit, viz. $T^{2}\propto a^{3}$, or $T\propto a^{3/2}$.(2) The cosmic factor $a$ of a spatially flat Einstein–de Sitter universe evolves as $a\propto t^{2/3}$, or $t\propto a^{3/2}$. These behaviors resemble the anisotropic time scaling, $\tau\propto l^{3/2}$, discussed in this paper.
@@
This predicted effect differs from the time dilation effect in GR, which is associated with the $g_{00}$ component of the spacetime metric [18, 19]. The new effect arises from the dependence of the clock rate on the dilaton field $\chi$, per Eq. (30). We investigate this new distinct phenomenon further elsewhere.  [^3] In addition to the time dilation effect in GR, there is a well-known effect in Special Relativity where two twice-intersecting time-like paths can have different total amounts of proper time in between. Therefore, we refer to our predicted time dilation effect as “the Third kind”.
Revisiting the Fujii–Wetterich scheme—We must emphasize that the dependence of clock rates on $\chi$ has been documented in the works of Fujii and Wetterich [4, 5, 6, 7, 8, 9], although it was not a focal point in their analysis. However, their findings significantly diverge from ours. Specifically, they reported the following relation for the clock rate  [^4] It is illuminating to quote Fujii [4]: “… the time and length in the microscopic unit frame are measured in units of $m^{-1}(t)$, in agreement with the physical situation that the time scale of atomic clocks, for example, is provided by the atomic levels which are determined by the Rydberg constant $(me^{4})^{-1}$ and Wetterich [7]: “The clock provided by the Hubble expansion in the standard description is now replaced by a clock associated to the increasing value of $\chi$.
@@
Discussion—Scalar degrees of freedom naturally emerge in various theoretical frameworks, including Kaluza–Klein, string theory, and braneworld scenarios [20]. By allowing matter to couple non-minimally with gravity via a scalar field (e.g., $\chi$), scalar–tensor theories can acquire scale symmetry, enabling them to evade observational constraints related to the fifth force [2, 3]. On the ground of dimensionality, it can be deduced that the lengthscale $l$ of a given physical process is determined by the value of $\chi$, per $l\propto\chi^{-1}$. This relationship justifies the use of the term “dilaton” for the field $\chi$.
In this paper, we imposed two requirements: (i) A dilaton field $\chi$ exists and directly couples with matter; and (ii) The intrinsic properties of matter—i.e., its gauge charge and inertial mass—are independent of the dilaton. Using the Lagrangian $\mathcal{L}_{\text{QED}}^{(\chi)}$ as a prototype, we demonstrated that these two prerequisites unambiguously lead to the following dependencies of $\hbar$ and $c$ on $\chi$: $\hbar_{\chi}\propto\chi^{-1/2}$ and $c_{\chi}\propto\chi^{1/2}$.
Based on these relationships, we then established that the timescale $\tau$ of a given physical process is related to the value of $\chi$ as $\tau\propto\chi^{-3/2}$. Together with $l\propto\chi^{-1}$, this leads to a universal anisotropic scaling law $\tau\propto l^{3/2}$ between the timescale and lengthscale of a physical process, resulting in a new time dilation effect. (Note: notwithstanding the resemblance of our anisotropic timescaling law to that in the Hořava–Lifshitz gravity [21], Lorentz symmetry is not broken in our approach.)
Three remarks are in order. Firstly, although $h_{\chi}$ and $c_{\chi}$ vary alongside $\chi$ in spacetime, the product $\hbar_{\chi}c_{\chi}$ remains constant. Secondly, in place of the QED prototype considered in this paper, an extension that incorporates the Higgs field and spontaneous symmetry breaking was presented in Ref. [22]. Its generalization to cover the Glashow–Weinberg–Salam model operating under electroweak symmetry breaking is also straightforward.
@@
Physical consequences—A varying $c$ in spacetime leads to direct and significant ramifications in theoretical and observational cosmology. Specifically, a variation in $c$ affects the propagation of light rays in an expanding universe, thereby altering standard cosmography both qualitatively and quantitatively. In [22, 23], we consider a VSL cosmology that relates the speed of light to the cosmic scale factor as $c\propto a^{-1/2}$. For this cosmology, we find that the canonical Lemaître redshift formula,$1+z=a^{-1}$, is no longer applicable and should be modified to $1+z=a^{-3/2}$. The $3/2$–exponent in this new formula arises from the anisotropic time scaling discussed in this paper. Consequently, our new Lemaître redshift formula necessitates a reanalysis for the Hubble diagram of Type Ia supernovae in the context of late-time accelerating expansion. In [22, 23], we conduct this reanalysis for the Pantheon Catalog using our VSL cosmology instead of the concordance $\Lambda$CDM model, and provide a new robust explanation for late-time cosmic acceleration based on VSL, while bypassing the need for dark energy.
@@ Acknowledgements.
I thank Cliff Burgess, Tiberiu Harko, Robert Mann, and Anne-Christine Davis for their constructive and supportive comments during the development of this work.

### v3

@@ Abstract
We reveal a novel aspect of scale-invariant actions that allow matter to couple with a dilaton field:The dynamics of the dilaton can induce variations in the Planck constant $\hbar$ and speed of light $c$.Our mechanism for generating variable $\hbar$ and $c$ in curved spacetimes via the dilaton offers a viablealternative account for late-time cosmic acceleration, bypassing the need for dark energy.
Motivation—Let us revisit the quantum electrodynamics (QED) of a spinor field $\psi$ with inertial mass $m$ and charge $e$, coupled with a $U(1)$ gauge vector field $A_{\mu}$ and embedded in a 4-dimensional spacetime described by the Einstein–Hilbert (EH) action of General Relativity (GR). With $\hbar$ and $c$ made explicit, the full action reads
@@
However, this relationship between $G$ and $\phi$ could cease to hold if matter couples non-minimally with gravity via the BD field. In this scenario, the involvement of $\phi$ in the matter Lagrangian, in principle, can cause both $c$ and $\hbar$ to vary alongside $\phi$ in spacetime, representing a departure from the standard BD paradigm. The current Letter explores this potentially far-reaching scenario.
@@
In the literature, when scale symmetry, also known as dilatation symmetry, is broken—specifically, when $\chi$ spontaneously acquires a non-vanishing vacuum expectation value—the process gives rise to a Goldstone mode conventionally referred to as a dilaton. We will adopt the terminology of “dilaton” for $\chi$ in this Letter. We should note, however, that the field $\chi$ in our action $\mathcal{S}_{2}$ can vary in spacetime, hence forfeiting its translational invariance.
Interestingly, a scale-invariant action of gravity and matter, such as $\mathcal{S}_{2}$, is able to evade observational bounds on the fifth force, a result established in [2, 3]. One can also add a “potential” term $V(\chi)$ to $\mathcal{L}_{\chi}$ provided that the added term respects scale symmetry. For example, $V(\chi)$ may contain terms like $\chi^{4}$ and $\chi^{-4}(\nabla\chi)^{4}$. Nevertheless, the inclusion of a scale-invariant potential will not affect the consideration presented in the rest of this Letter.
@@
$G=\frac{\hat{c}^{3}}{16\pi\hat{\chi}^{2}\hat{\hbar}}$ (15)
Importantly, per Eq. (14), the charge $e$ and inertial mass $m$ are independent of $\chi$. This result is desirable, as both charge and inertial mass—being intrinsic properties of a particle—should be oblivious to external factors, in particular the background dilaton field $\chi$. Befittingly, in our scheme, rather than making mass variable, the effect of the dilaton is translated into $\hbar_{\chi}$ and $c_{\chi}$, which respectively regulate the quantization and propagation of fields against a background spacetime.
@@
Moreover, our scheme brings about an additional benefit: not only is the $U(1)$ gauge charge $e$—which specifies the strength of electromagnetic interaction—independent of $\chi$, but the Newton gravitational constant $G$—characterizing the strength of gravitational interaction—is also independent of $\chi$, as evident in Eq. (15). Note that in our scheme, the Planck mass, given by $m_{\chi}^{\text{Planck}}\,:=$$\sqrt{\frac{\hbar_{\chi}c_{\chi}}{G}}=$$\,\sqrt{16\pi}\,\frac{\hat{\chi}\hat{\hbar}}{\hat{c}}$, is constant rather than variable.  [^1] Our Letter does not concern with the hierarchy issue, namely why $m_{\chi}^{\text{Planck}}\gg m$, or equivalently why $\mu\ll\sqrt{16\pi}$.
Two comments are necessary. Firstly, the dimensionless parameters $\alpha$ and $\mu$ are “running coupling constants” in the renormalization group flow of $\mathcal{L}_{\text{QED}}^{(\chi)}$ when loop corrections involving $\psi$ and $A_{\mu}$ are included. Thus, although $e$ and $m$ are independent of $\chi$, they can “run” as functions of the momentum scale at which they are measured.
Secondly, per Eq. (10), the fine structure constant $\alpha$ is unrelated to $\chi$ and thus does not vary in spacetime. This criterion is compatible with the theoretical bound that time-variations in $\alpha$, if caused by a cosmic scalar field, would be of order $\left|\frac{\delta\,\alpha}{\alpha}\right|<10^{-37}$ since the earliest stages of galaxy formation [20]. (This bound hence casts doubt on the validity of the observation reported in [21].)
@@
Within the open set, $\mathcal{L}_{\text{QED}}^{\text{ eff}}$ describes the quantum electrodynamics of a spinor field with charge $e$ and inertial mass $m$, specified in Eq. (14), coupled with a $U(1)$ gauge vector field. It is important to emphasize that, unlike the original $\mathcal{L}_{\text{QED}}$ in Eq. (3), the quantization and propagation of fields in $\mathcal{L}_{\text{QED}}^{\text{ eff}}$ are governed by the effective quantum of action $\hbar_{\chi}$ and speed of light $c_{\chi}$, which in turn are determined by the dilaton $\chi$ via Eq. (13). More generally, as $\chi$ varies across the manifold, different locations then correspond to separate replicas of the effective Lagrangian $\mathcal{L}_{\text{QED}}^{\text{ eff}}$, each operating with its respective values of $\hbar_{\chi}$ and $c_{\chi}$. Conceptually, the dynamics of $\chi$ thus induces variations in $\hbar$ and $c$ across spacetime.  [^2] The co-variability of $\hbar$ and $c$ has been considered in Refs. [23, 25, 24, 22], although a concrete mechanism for generating their variations in spacetime is still lacking in these works.
We emphasize that although the action described in Eqs. (6), (7) and (8) has been previously explored in both the contexts of gravitation and particle physics (e.g., in Refs. [3, 4, 5, 6, 7, 8, 9, 2, 17, 14, 12, 19, 11, 13, 16, 10, 15, 18]), its capacity to produce variations in $\hbar$ and $c$ in spacetime—as presented in this Letter thus far—has not been recognized or documented.
@@
The action $\mathcal{S}_{2}$ is generally covariant, whereas the action associated with $\mathcal{L}_{\text{QED}}^{\text{ eff}}$ is locally Lorentz invariant. Our approach is therefore compatible with the varying–$c$ framework that respects both general covariance and local Lorentz invariance, as laid out by Magueijo in [32].
@@
$i\hbar_{\chi}\,\frac{\partial}{\partial t}\bigl{|}\Psi\bigr{\rangle}=\hat{H}\bigl{|}\Psi\bigr{\rangle}$ (17)
A physical consequence is that, for two locations with different dilaton values $\chi_{1}$ and $\chi_{2}$, the evolution of a quantum state occurs at different rates determined by each respective $h_{\chi}$ value. The discrepancy in the clock rates at the two locations gives rise to a new time dilation effect, resulting from the variation in $\hbar$ induced by the dynamics of $\chi$. We will briefly discuss this concrete definitive prediction in a later section of the Letter.
@@
The energy level of an (relativistic) electron in a quantum state $\left|n,\,j\right\rangle$ of a Hydrogen atom is a well-established result; it is given by [33]  [^3] Beyond the Dirac theory, the two states ${}^{2}S_{1/2}$ and ${}^{2}P_{1/2}$ of the Hydrogen atom become non-degenerate due to radiative corrections, a physical effect known as the Lamb shift. Their energy difference is given by $\frac{1}{6\pi}\alpha^{5}mc^{2}\ln\frac{1}{\pi\alpha}$ [34] which is also proportional to $\chi$ due to $c\propto\chi^{1/2}$. We thank the Referee for suggesting this connection.
@@
$A=\frac{4\alpha}{3}\frac{m^{3}c_{\chi}^{4}}{h_{\chi}^{3}}\left\langle\vec{r}_{if}\right\rangle^{2}$ (26)
@@
This result is one of the key findings of our Letter.  [^4] There are two examples in classical physics that carry a $3/2$–exponent hallmark: (1) Kepler’s Third law, which states that the square of a planet’s orbital period $T$ is proportional to the cube of the semi-major axis $a$ of its orbit, viz. $T^{2}\propto a^{3}$, or $T\propto a^{3/2}$.(2) The cosmic factor $a$ of a spatially flat Einstein–de Sitter universe evolves as $a\propto t^{2/3}$, or $t\propto a^{3/2}$. These behaviors resemble the anisotropic time scaling, $\tau\propto l^{3/2}$, discussed in this Letter.
@@
This predicted effect differs from the time dilation effect in GR which is associated with the $g_{00}$ component of the spacetime metric [36, 37]. It arises from the dependence of the clock rate on the dilaton field $\chi$, per Eq. (30).  [^5] In addition to the time dilation effect in GR, there is a well-known effect in Special Relativity where two twice-intersecting time-like paths can have different total amounts of proper time in between. Therefore, we refer to our predicted time dilation effect as “the Third kind”. We emphasize that this new effect is physical, meaning it is, in principle, measurable by comparing the time lapses of two clocks situated at two separate locations with different values of the dilaton field. We investigate this distinct phenomenon further elsewhere.
Revisiting the Fujii–Wetterich scheme—We must emphasize that the dependence of clock rates on $\chi$ has been documented in the works of Fujii and Wetterich [4, 5, 6, 7, 8, 9], although it was not a focal point in their analysis. However, their findings significantly diverge from ours. Specifically, they reported the following relation for the clock rate  [^6] We find it illuminating to quote Fujii [4]: “… the time and length in the microscopic unit frame are measured in units of $m^{-1}(t)$, in agreement with the physical situation that the time scale of atomic clocks, for example, is provided by the atomic levels which are determined by the Rydberg constant $(me^{4})^{-1}$ and Wetterich [7]: “The clock provided by the Hubble expansion in the standard description is now replaced by a clock associated to the increasing value of $\chi$.
@@
Discussions—Scalar degrees of freedom naturally emerge in various theoretical frameworks, including Kaluza–Klein, string theory, and braneworld scenarios [38]. By allowing matter to couple non-minimally with gravity via a scalar field (e.g., $\chi$), scalar–tensor theories can acquire scale symmetry, enabling them to evade observational constraints related to the fifth force [2, 3]. On the ground of dimensionality, it can be deduced that the lengthscale $l$ of a given physical process is determined by the value of $\chi$, per $l\propto\chi^{-1}$. This relationship justifies the use of the term “dilaton” for the field $\chi$.
In this Letter, we imposed two requirements: (i) A dilaton field $\chi$ exists and directly couples with matter; and (ii) The intrinsic properties of matter—i.e., its gauge charge and inertial mass—are independent of the dilaton. Using the Lagrangian $\mathcal{L}_{\text{QED}}^{(\chi)}$ as a prototype, we demonstrated that these two criteria unambiguously lead to the following dependencies of $\hbar$ and $c$ on $\chi$: $\hbar_{\chi}\propto\chi^{-1/2}$ and $c_{\chi}\propto\chi^{1/2}$.
Based on these relationships, we then established that the timescale $\tau$ of a given physical process is related to the value of $\chi$ as $\tau\propto\chi^{-3/2}$. Together with $l\propto\chi^{-1}$, this leads to a universal anisotropic scaling law $\tau\propto l^{3/2}$ between the timescale and lengthscale of a physical process, resulting in a new time dilation effect. (Note: notwithstanding the resemblance of our anisotropic timescaling law to that in Hořava–Lifshitz gravity [39], Lorentz symmetry is not broken in our approach which respects both general covariance and local Lorentz invariance.)
Three remarks are in order. Firstly, although $h_{\chi}$ and $c_{\chi}$ vary alongside $\chi$ in spacetime, the product $\hbar_{\chi}c_{\chi}$ remains constant. Secondly, in place of the QED prototype considered in this Letter, an extension that incorporates the Higgs field and spontaneous symmetry breaking was presented in Ref. [40]. Its generalization to cover the Glashow–Weinberg–Salam model operating under electroweak symmetry is also straightforward.
@@
Physical consequences—A varying $c$ in spacetime leads to direct and significant ramifications in theoretical and observational cosmology. Specifically, a variation in $c$ affects the propagation of light rays in an expanding universe, thereby altering standard cosmography both qualitatively and quantitatively. In [41], we consider a VSL cosmology that relates the speed of light to the cosmic scale factor as $c\propto a^{-1/2}$. For this cosmology, we find that the canonical Lemaître redshift formula, $1+z=a^{-1}$, is no longer applicable and should be modified to $1+z=a^{-3/2}$. The $3/2$–exponent in this new formula arises from the anisotropic time scaling discussed in this Letter. Consequently, our new Lemaître redshift formula necessitates a reanalysis for the Hubble diagram of Type Ia supernovae in the context of late-time accelerating expansion. In [41], we conduct this reanalysis for the Pantheon Catalog using our VSL cosmology instead of the concordance $\Lambda$CDM model, and provide a new robust explanation for late-time cosmic acceleration based on VSL, while bypassing the need for dark energy.
It is important to note that the observational bounds established in the literature in support of a constant speed of light have predominantly relied on standard cosmology [47, 48, 52, 53, 57, 43, 44, 45, 46, 49, 50, 51, 56, 42, 55, 54, 58]. However, our new Lemaître redshift formula represents a critical departure from this framework, thereby warranting a reevaluation of these constraints. The consensus regarding the absence of variation in $c$ in observational cosmology must be reconsidered in light of our findings, prompting a comprehensive reanalysis of these established bounds.
@@
Conclusion—Symmetries play an instrumental role in constructing modern physical theories. Prominent examples are general covariance in GR and gauge invariance in the Glashow–Weinberg–Salam model of particle physics. In this Letter, we examined a scale-invariant action specified by Eqs. (6), (7) and (8), which is generally covariant, locally Lorentz invariant, and $U(1)$ gauge invariant (for the matter sector). The scale invariance necessitates a non-minimal coupling of matter to gravity via a dilaton. Although this theory has been extensively investigated [3, 4, 5, 6, 7, 8, 9, 2, 17, 14, 12, 19, 11, 13, 16, 10, 15, 18], its capacity to permit variable Planck constant and speed of light in curved spacetimes has been overlooked in the literature. We addressed this shortcoming by providing—for the first time—a concrete mechanism to generate variations in $\hbar$ and $c$ across spacetimes, leading to material consequences in both theoretical and observational cosmology, as presented in Ref. [41].
@@ Acknowledgements.
I thank Clifford Burgess, Tiberiu Harko, Robert Mann and Anne-Christine Davis for their constructive and supportive comments during the development of this work. I thank the anonymous Referee for his or her significant feedback in improving this Letter.